OpenStudy (anonymous):

solve x^2 + 8x + 16 = 0

Directrix (directrix):

(x + 4) (x + 4) = 0 Zero Product Property x + 4 = 0 or x + 4 = 0 x = -4 or x = -4 -4 is a double root.

Directrix (directrix):

Take a look at the graph: http://www.wolframalpha.com/input/?i=x%5E2+%2B+8x+%2B+16+%3D+0